OpenStudy (anonymous):

6-3/20 X 10 6-3/20 X 10 @Mathematics

OpenStudy (diyadiya):

3/20 x 10 30/20 =3/2

OpenStudy (anonymous):

u still have to minus 6 right

OpenStudy (diyadiya):

is it \[6- \frac{3}{20}\] or \[\frac{6-3}{20}\]